1. A 751 mL NaCl solution is diluted to a volume of 1.06 L and a concentration of 8.00 M . What was the initial concentration?

Express your answer with the appropriate units.

2. How many liters of a 3.18  M K2SO4 solution are needed to provide 98.8  g of K2SO4 (molar mass 174.01 g/mol)? Recall that M is equivalent to mol/L.

Express your answer to three significant figures.
Answer #1

1)

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M2 = 8 M

V1 = 751 mL

V2 = 1060 mL

use:

M1*V1 = M2*V2

M1 = (M2 * V2) / V1

M1 = (8*1060)/751

M1 = 11.3 M

Answer: 11.3 M

Only 1 question at a time please